# CoVR: Learning Composed Video Retrieval from Web Video Captions

### v3

@@ Appendix B Implementation details
We describe the dataset generation computation time (Section B.1), further training details (Section B.2), provide the templates we use for our rule-based baseline (Section B.3), and details about our MTG-LLM finetuning and inference (Section B.4).
B.1 Dataset generation computation time
@@
  • Caption similarity search: To identify captions with one-word differences, we employed the faiss library (Johnson, Douze, and Jégou 2019) to select the 100 closest captions, avoiding the need to compare each caption against the entire set of 2 million captions. This optimization significantly reduced the search time, resulting in 2.5 hours.
@@
B.2 Training details
Here, we provide implementation details in addition to Section 4.1 of the main paper. In terms of the optimization algorithm, we utilize AdamW (Loshchilov and Hutter 2019). For our MTG-LLM, we finetune for one epoch with a batch size of 128 and a learning of $3\mathrm{e}{-5}$ that is warmed up linearly for the first $100$ steps and then kept constant. For our CoVR model, keeping the visual backbone frozen largely improves the efficiency of the training process: an epoch on the CIRR dataset takes 4 minutes with a frozen backbone and 25 minutes with a finetuned backbone, while leading to similar performance. During the training process, we employ several image data augmentations. These transformations include a random resized crop, where the input image is resized to a resolution of $384\times 384$. Additionally, we apply a random horizontal flip and random adjustments to contrast, brightness, sharpness, translation, and rotation. We use a weight decay of 0.05 and an initial learning rate of $1\mathrm{e}{-5}$ that is decayed to 0 following a cosine schedule over 10 epochs.
B.3 List of rule-based templates
In the ablation studies (Section 4.4 of the main paper), we introduced a rule-based MTG baseline. Here, in Table A.1, we show the templates used for the rules. We refer to Section D.2 (Table A.11) for qualitative comparison with our finetuned MTG-LLM.
@@ B.3 List of rule-based templates
B.4 Generating a modification text from paired captions with MTG-LLM
@@
As mentioned in Section 3.1 of the main paper, we add 15 manually prepared text triplets to the existing 700 text triplets from (Brooks, Holynski, and Efros 2023) used for training. The motivation is to address specific CoVR cases not present in the original set of triplets, such as “remove clouds and reveal only sky” given input captions “Clouds timelapse” and “Sky timelapse”. We show these 15 samples in Table A.2.
Table A.2: Added examples to the MTG-LLM training: We add the below 15 examples to the set of 700 text triplets from (Brooks, Holynski, and Efros 2023).
@@
Appendix C Additional experiments
We provide additional experiments, reporting CoVR results obtained by training on data generated with prompting (i.e., without finetuning) the LLM (Section C.1), results when changing the visual query from an image to a video (Section C.2), varying the pretrained BLIP model (Section C.3), incorporating visual similarity between videos (Section C.4), results when filtering dynamic or static videos (Section C.5) effect of the modification text length (Section C.6), optimal number of frames (Section C.7), and full CoIR metrics (Section C.8).
C.1 Prompting versus finetuning the MTG-LLM
Here, we justify why we finetuned Llama as opposed to simply prompting it without any training. For prompting, we prepend few-shot examples of pairs of captions and desired generated texts, before adding the two captions in question. In particular, we use the following sentence:
@@ C.1 Prompting versus finetuning the MTG-LLM
Then, we concatenate our two captions for which we wish to generate a modification text. Table A.3 shows that finetuning the MTG-LLM for generating the training data is much more effective than prompting it without finetuning, as measured by CoVR performance on WebVid-CoVR-Test and CoIR performance on CIRR. This is also consistent with our qualitative observations: we found that the LLM struggles to perform the modification text generation without finetuning (see Table A.11 in the next section).
Table A.3: Prompting versus finetuning LLM: We compare our finetuned model (MTG-LLM) to a prompting baseline (see Section C.1) and observe important gains in the downstream performance of the model trained on the generated data.
WebVid-CoVR-Test CIRR
Model R@1 R@5 R@10 R@50 R@1 R@5 R@10 R@50
Prompting 51.33 76.68 85.13 96.71 35.59 64.67 76.19 92.87
Finetuning 53.13 79.93 86.85 97.69 39.28 68.22 78.94 94.65
C.2 Video query for CoVR
As noted in Section 3 of the main paper, we focus on image queries in this paper. This was because querying with an image has arguably more applications for realistic search scenarios. Here, we explore the setup of using a video as the visual query instead of an image query. We can do this since our dataset consists of video-text-video triplets. To encode a query video, we sample 5 equally-spaced frames and compute visual embeddings for each frame using the BLIP image encoder. We then average the per-frame embeddings and forward it through the BLIP cross-attention layers to obtain a multimodal query embedding $f(q,t)$. Note that we keep the target video representation fixed to 15 frames with weighted embedding averaging as described in Section 3.3 of the main paper. As seen in Table A.4, using 5 query frames leads to similar performance to using the middle frame.
Table A.4: Querying with a video: We report results on WebVid-CoVR-Test by using multiple frames from the query video. Recall that the rest of the paper investigates the setup where the middle video frame is used as an image query. To keep the computational complexity low, we only use 5 query video frames (uniformly sampled throughout the video). The number of target video frames remains unchanged as 15. The performance is similar to the image query setup, with marginal increase.
@@ C.2 Video query for CoVR
Image (middle frame) 53.17 79.93 86.85 97.69
Video (5 uniform frames) 53.91 79.85 87.09 97.42
C.3 Variants of pretrained BLIP backbones
All experiments in this paper are performed with the BLIP model (Li et al. 2022b) finetuned on COCO (Lin et al. 2014). Here, we include experiments when changing this backbone with other pretrained BLIP variants. Specifically, we use the BLIP model without COCO finetuning (BLIP base), and the BLIP model finetuned on Flickr30k (Plummer et al. 2015). For this experiment (as in the last row of Table 2 of the main paper), we use pretrained cross-attention layers of BLIP as our multimodal combined representation, and finetune them on WebVid-CoVR with 15-frame target video embeddings. In Table A.5, we observe that the BLIP model finetuned on COCO has the highest performance.
Table A.5: Variants of pretrained BLIP backbones: We compare the BLIP model without finetuning (base), BLIP finetuned on Flickr30k, and BLIP finetuned on COCO (the one used in the rest of the paper) (Li et al. 2022b). For this experiment, we finetune the models on WebVid-CoVR using the cross-attention layers of BLIP as the fusion method, and 15 frames for the target video as in the last row of Table 2.
@@ C.3 Variants of pretrained BLIP backbones
BLIP Base 50.74 78.91 86.23 97.34
BLIP Flickr30k 52.50 79.46 86.70 97.77
BLIP COCO 53.13 79.93 86.85 97.69
C.4 Incorporating visual similarity between videos
@@
C.5 Dynamic vs static content
The videos in our WebVid-CoVR dataset contain both dynamic sequences with motion, as well as static content. To analyze the prevalence, we compute the optical flow using the Gunnar Farneback’s algorithm (Farnebäck 2003) and empirically choose a magnitude threshold of 1 to distinguish between videos with static and dynamic elements. The magnitude value is obtained by averaging the Euclidean norms of motion vectors in both horizontal and vertical directions across the computed video frames. We identify that around 25% of the triplets contain static target videos, which represents approximately 21% of the overall target videos.
As a further experiment, we train models omitting either only the static portion or only the dynamic portions. The results in Table A.7 show that training on both dynamic and static triplets is beneficial, with a minor decrease in performance if we omit these static videos during training (while maintaining the same iteration count). This may be because image training data can still be complementary to video training (Bain et al. 2021).
@@ C.5 Dynamic vs static content
Table A.7: Performance on WebVid-CoVR-Test when training on dynamic vs static video triplets: We see that training on both static and dynamic videos is beneficial.
@@
Static 25% 49.26 76.41 84.47 97.03
Dynamic 75% 52.39 79.38 85.49 97.42
Static + Dynamic 100% 53.13 79.93 86.85 97.69
C.6 Effect of modification text length
@@
(ours) 23.36 53.13 79.93 39.06 70.33 27.70 44.63
(longest) 33.35 52.54 78.72 36.67 68.48 24.55 40.49
@@
This shows that while the generated modification texts are not as long on average as the ones from CIRR, our generated texts still provide useful training signal, as evidenced by the state-of-the-art performance of models trained on our dataset when transferred to standard CoIR tasks (CIRR and FashionIQ in Section A.9 and Table A.10). Moreover, the average number of characters of the modification text in WebVid-CoVR and the widely used FashionIQ CoIR benchmark are comparable (see Table 1 of the main paper: 23.36 vs 27.13, respectively).
Mode Pretraining Recall@K $\text{R}_{\text{subset}}$@K
Method Data K=1 K=5 K=10 K=50 K=1 K=2 K=3
Train (CIRR) TIRG (Vo ’19$\dagger$ - 14.61 48.37 64.08 90.03 22.67 44.97 65.14
TIRG+LastConv (Vo ’19) $\dagger$ - 11.04 35.68 51.27 83.29 23.82 45.65 64.55
MAAF (Doods ‘20$\dagger$ - 10.31 33.03 48.30 80.06 21.05 41.81 61.60
MAAF-BERT (Doods ‘20$\dagger$ - 10.12 33.10 48.01 80.57 22.04 42.41 62.14
MAAF-IT (Doods ‘20$\dagger$ - 09.90 32.86 48.83 80.27 21.17 42.04 60.91
MAAF-RP (Doods ‘20$\dagger$ - 10.22 33.32 48.68 81.84 21.41 42.17 61.60
ARTEMIS (Delmas ‘22) - 16.96 46.10 61.31 87.73 39.99 62.20 75.67
CIRPLANT (Liu ‘21$\dagger$ - 19.55 52.55 68.39 92.38 39.20 63.03 79.49
LF-BLIP (Levy 2023) - 20.89 48.07 61.16 83.71 50.22 73.16 86.82
CompoDiff (Gu ‘23) ST18M 22.35 54.36 73.41 91.77 35.84 56.11 76.60
Combiner (Baldrati ‘22) - 33.59 65.35 77.35 95.21 62.39 81.81 92.02
CASE (Levy ‘23) - 48.00 79.11 87.25 97.57 75.88 90.58 96.00
CASE (Levy ‘23) LaSCo 48.68 79.98 88.51 97.49 76.39 90.12 95.86
CASE (Levy ‘23) LaSCo+COCO 49.35 80.02 88.75 97.47 76.48 90.37 95.71
CoVR-BLIP - 49.16 79.76 88.65 97.49 78.12 91.25 96.26
CoVR-BLIP WebVid-CoVR 50.41 80.96 89.35 97.64 77.25 91.04 96.34
[…]
CompoDiff (Gu ‘23) ST18M 19.37 53.81 72.02 90.85 28.96 49.21 67.03
Pic2Word (Saito ‘23) CC3M 23.90 51.70 65.30 87.80 - - -
CASE (Levy ‘23) LaSCo 30.89 60.75 73.88 92.84 60.17 80.17 90.41
CASE (Levy ‘23) LaSCo+COCO 35.40 65.78 78.53 94.63 64.29 82.66 91.61
CoVR-BLIP - 19.95 41.98 52.31 73.54 64.72 83.42 91.95
CoVR-BLIP WebVid-CoVR 39.28 68.22 78.94 94.65 71.23 86.22 94.02
Table A.9: State-of-the-art comparison on the CIRR test set: In Table 3 of the main paper, we presented a subset of the metrics on CIRR dataset due to space constraints. Here, we provide the full set of metrics. Specifically, we add three columns for each $Recall_{\text{subset}}$ metric in CIRR. $\dagger$ denotes results reported by (Liu et al. 2021). For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. CC3M denotes Conceptual Captions 3M (Sharma et al. 2018), ST18M denotes  (Gu et al. 2023), and COCO denotes (Lin et al. 2014).
Mode Pretraining Dress Shirt Toptee Average
Method Data R@10 R@50 R@10 R@50 R@10 R@50 R@10 R@50
Train (FashionIQ) JVSM (Chen ‘20) - 10.70 25.90 12.00 27.10 13.00 26.90 11.90 26.60
CIRPLANT (Liu ‘21) - 17.45 40.41 17.53 38.81 21.64 45.38 18.87 41.53
TRACE w/BER (Jandial ‘20) - 22.70 44.91 20.80 40.80 24.22 49.80 22.57 46.19
VAL w/GloVe (Chen ‘20) - 22.53 44.00 22.38 44.15 27.53 51.68 24.15 46.61
MAAF (Dodds ‘20) - 23.80 48.60 21.30 44.20 27.90 53.60 24.30 48.80
CurlingNet (Yu ‘20$\ddagger$ - 26.15 53.24 21.45 44.56 30.12 55.23 25.90 51.01
RTIC-GCN (Shin ‘21$\ddagger$ - 29.15 54.04 23.79 47.25 31.61 57.98 28.18 53.09
CoSMo (Lee ‘21) - 25.64 50.30 24.90 49.18 29.21 57.46 26.58 52.31
ARTEMIS (Delmas ‘22) - 27.16 52.40 21.78 43.64 29.20 53.83 26.05 50.29
DCNet (Kim ‘21) - 28.95 56.07 23.95 47.30 30.44 58.29 27.78 53.89
SAC w/BERT (Jandial ‘22) - 26.52 51.01 28.02 51.86 32.70 61.23 29.08 54.70
FashionVLP (Goenka ‘22) - 32.42 60.29 31.89 58.44 38.51 68.79 34.27 62.51
LF-CLIP (Baldrati ‘22) - 31.63 56.67 36.36 58.00 38.19 62.42 35.39 59.03
LF-BLIP (Levy ‘23) - 25.31 44.05 25.39 43.57 26.54 44.48 25.75 43.98
CASE (Levy ‘23) LaSCo 47.44 69.36 48.48 70.23 50.18 72.24 48.79 70.68
CoVR-BLIP - 43.51 67.94 48.28 66.68 51.53 73.60 49.02 70.24
CoVR-BLIP WebVid-CoVR 44.55 69.03 48.43 67.42 52.60 74.31 49.40 70.98
ZS Random - 00.26 01.31 00.16 00.79 00.19 00.95 00.06 00.32
Pic2Word (Saito et al. 2023) CC3M 20.00 40.20 26.20 43.60 27.90 47.40 24.70 43.70
CoVR-BLIP - 13.48 31.96 16.68 30.67 17.84 35.68 16.00 32.77
CoVR-BLIP WebVid-CoVR 21.95 39.05 30.37 46.12 30.78 48.73 27.70 44.63
Table A.10: State-of-the-art comparison on the FashionIQ validation set: In Table 3 of the main paper, we presented a subset of the metrics on FashionIQ dataset due to space constraints. Here, we provide the full set of metrics. Specifically, we add individual categories (Dress, Shirt, and Toptee). For methods that pretrain specifically for composed retrieval, we indicate their pretraining data. $\ddagger$ denotes results reported by (Shin et al. 2021). CC3M denotes Conceptual Captions 3M (Sharma et al. 2018). ZS stands for Zero Shot.
C.7 Optimal number of frames
In Section 4.2 we saw that training with 15 frames further boosts the performance over using only the middle frame. We provide further analysis by incrementally increasing the number of frames for training and testing. We report the average recall on WebVid-CoVR-Test and observe a steady increase in performance with more frames (see Figure A.8).
C.8 Full CoIR metrics
In Table 3 of the main paper, we presented a subset of the metrics on FashionIQ and CIRR datasets due to space constraints. Here, we provide the full set of metrics in Table 3 on CIRR and Table A.9 on FashionIQ. Specifically, we add three columns for each $Recall_{\text{subset}}$ metric in CIRR. For FashionIQ, we add individual categories (Dress, Shirt, and Toptee), as well as the aggregated average results across these categories.
Appendix D Qualitative analysis

### v4

@@ C Implementation details
We describe the dataset generation computation time (Section C.1), further training details (Section C.2), provide the templates we use for our rule-based baseline (Section C.3), details about our MTG-LLM finetuning and inference (Section C.4), and prompt to our prompting experiment (Section C.5.
C.1 Dataset generation computation time
@@
  • Caption similarity search: To identify captions with one-word differences, we employed the faiss library [76] to select the 100 closest captions, avoiding the need to compare each caption against the entire set of 2 million captions. This optimization significantly reduced the search time, resulting in 2.5 hours.
@@
C.2 Training details
Here, we provide implementation details in addition to Section 4.1 of the main paper. In terms of the optimization algorithm, we utilize AdamW [77]. For our MTG-LLM, we finetune for one epoch with a batch size of 128 and a learning of $3\mathrm{e}{-5}$ that is warmed up linearly for the first $100$ steps and then kept constant. For our CoVR model, keeping the visual backbone frozen largely improves the efficiency of the training process: an epoch on the CIRR dataset takes 4 minutes with a frozen backbone and 25 minutes with a finetuned backbone, while leading to similar performance. During the training process, we employ several image data augmentations. These transformations include a random resized crop, where the input image is resized to a resolution of $384\times 384$. Additionally, we apply a random horizontal flip and random adjustments to contrast, brightness, sharpness, translation, and rotation. We use a weight decay of 0.05 and an initial learning rate of $1\mathrm{e}{-5}$ that is decayed to 0 following a cosine schedule over 10 epochs.
C.3 List of rule-based templates
In the ablation studies (Section 4.6 of the main paper), we introduced a rule-based MTG baseline. Here, in Table A.3, we show the templates used for the rules. We refer to Section E.2 (Table A.11) for qualitative comparison with our finetuned MTG-LLM.
@@ C.3 List of rule-based templates
C.4 Generating a modification text from paired captions with MTG-LLM
@@
As mentioned in Section 3.1 of the main paper, we add 15 manually prepared text triplets to the existing 700 text triplets from [11] used for training. The motivation is to address specific CoVR cases not present in the original set of triplets, such as “remove clouds and reveal only sky” given input captions “Clouds timelapse” and “Sky timelapse”. We show these 15 samples in Table A.4.
TABLE A.4: Added examples to the MTG-LLM training: We add the below 15 examples to the set of 700 text triplets from [11].
@@
C.5 Details of the LLaMA prompt
In the ablation studies (Section 4.6 of the main paper), we justified why we finetuned LLaMA as opposed to simply prompting it without any training. Here, we show how we determine the prompt for the aforementioned experiment. Specifically, we prepend few-shot examples of pairs of captions and desired generated texts, before adding the two captions in question. In particular, we use the following sentence:
@@ C.5 Details of the LLaMA prompt
Then, we concatenate our two captions for which we wish to generate a modification text. The previous results in (Table VIII of the main paper, are also consistent with our qualitative observations: we found that the LLM struggles to perform the modification text generation without finetuning (see Table A.11 in the next section).
D Additional experiments
We provide additional experiments, reporting CoVR results when changing the visual query from an image to a video (Section D.1), effect of backbones (Section D.2), incorporating visual similarity between videos (Section D.3), results when filtering dynamic or static videos (Section D.4) effect of the modification text length (Section D.5), and optimal number of frames (Section D.6).
D.1 Video query for CoVR
As noted in Section 3 of the main paper, we focus on image queries in this paper. This was because querying with an image has arguably more applications for realistic search scenarios. Here, we explore the setup of using a video as the visual query instead of an image query. We can do this since our dataset consists of video-text-video triplets. To encode a query video, we sample 15 equally-spaced frames and compute visual embeddings for each frame using the BLIP-2 image encoder. We then average the per-frame embeddings and forward it through the BLIP-2 cross-attention layers to obtain a multimodal query embedding $f(q,t)$. Note that we keep the target video representation fixed to 15 frames with weighted embedding averaging as described in Section 3.3 of the main paper. As seen in Table A.5, using 15 query frames leads to similar performance to using the middle frame.
TABLE A.5: Querying with a video: We report results on WebVid-CoVR-Test by using multiple frames from the query video. Recall that the rest of the paper investigates the setup where the middle video frame is used as an image query. We use 5 query video frames (uniformly sampled throughout the video). The number of target video frames remains unchanged as 15. The performance is similar to the image query setup, with marginal increase.
@@ D.1 Video query for CoVR
Image (middle frame) 59.82 83.84 91.28 98.24
Video 59.55 84.19 90.85 98.32
D.2 Effect of backbones
Variants of pretrained BLIP-2 models. All experiments in this paper are performed with the BLIP-2 model [16] finetuned on COCO [61]. Here, we include experiments when changing this backbone with ViT-L without COCO finetuning (BLIP-2 base). For this experiment (as in the last row of Table II of the main paper), we use pretrained cross-attention layers of BLIP-2 as our multimodal combined representation, and finetune them on WebVid-CoVR. In Table A.6, we observe that the BLIP-2 model fine-tuned with COCO has a similar performance to BLIP-2 Base, with a slight improvement on R@1.
TABLE A.6: Variants of pretrained BLIP-2 backbones: We compare the BLIP-2 model without finetuning (base) and BLIP-2 finetuned on COCO (the one used in the rest of the paper) [1]. For this experiment, we finetune the models on WebVid-CoVR using the cross-attention layers of BLIP-2 as the fusion method.
@@ D.2 Effect of backbones
BLIP-2 Base 59.66 84.04 90.92 98.32
BLIP-2 COCO 59.82 83.84 91.28 98.24
Effect on CIR benchmarks. As shown in Table II of the main paper, the BLIP-2 backbone with cross-attention layer finetuning achieves the highest performance on the WebVid-CoVR-Test dataset. In Table A.7, we extend the comparison to additional CIR benchmarks, evaluating frozen CLIP with average fusion (no training), CLIP with MLP fusion, and BLIP and BLIP-2 with cross-attention layer finetuning. The results show that MLP fusion fails to generalize effectively across these datasets when trained with WebVid-CoVR, while the BLIP-2 backbone continues to deliver the best results across all benchmarks.
TABLE A.7: CLIP vs BLIP performance on CIR benchmarks: The table compares frozen CLIP with average fusion, CLIP with MLP fusion, and BLIP and BLIP-2 with cross-attention layer finetuning. Results demonstrate that while MLP fusion struggles to generalize across datasets when trained with WebVid-CoVR, the BLIP-2 backbone consistently outperforms CLIP across all benchmarks.
Backbone Pretraining CIRR FashionIQ CIRCO
Data R@1 R@10 mAP@5
CLIP - 10.65 17.63 3.63
CLIP + MLP WV-CC-CoVIR 11.39 19.10 5.67
BLIP WV-CC-CoVIR 38.48 27.70 21.43
BLIP-2 WV-CC-CoVIR 43.74 38.15 28.29
D.3 Incorporating visual similarity between videos
@@
D.4 Dynamic vs static content
The videos in our WebVid-CoVR dataset contain both dynamic sequences with motion, as well as static content. To analyze the prevalence, we compute the optical flow using the Gunnar Farneback’s algorithm [78] and empirically choose a magnitude threshold of 1 to distinguish between videos with static and dynamic elements. The magnitude value is obtained by averaging the Euclidean norms of motion vectors in both horizontal and vertical directions across the computed video frames. We identify that around 25% of the triplets contain static target videos, which represents approximately 21% of the overall target videos.
As an additional method to distinguish between static and dynamic triplets, we analyze the part of speech of the word that changes in the target video caption. Specifically, we examine whether the modified word is a noun, verb, or another part of speech. We find that 65% of the changes involved nouns, 9% involve verbs, and the remaining 26% correspond to other parts of speech (such as adjectives).
We train models omitting either only the static portion or only the dynamic portions. The results in Table A.9 show that training on both dynamic and static triplets is beneficial, with a minor decrease in performance if we omit static videos during training (while maintaining the same iteration count). This may be because image training data can still be complementary to video training [12].
@@ D.4 Dynamic vs static content
Ideally, leveraging dynamic videos would imply capturing not just static states but also the transitions between them, which is a key distinction between composed video retrieval and composed image retrieval. For instance, when given an image of ’wheat flour’ and a text query like ’change this to dough,’ an image search will likely retrieve the end result, such as ’dough.’ In contrast, composed video retrieval could potentially capture the entire process, showing ’how’ the transformation from ’flour’ to ’dough’ occurs, providing a richer and more informative result. We leave this potential application to future research, which will benefit from better datasets and improved modeling techniques.
TABLE A.9: Performance on WebVid-CoVR-Test when training on dynamic vs static video triplets: We employ two methods to classify videos as static or dynamic: one based on the optical flow of the target video, and another based on the type of word (noun or verb) modified in the target caption. The results indicate that training on the full dataset, which includes both videos with temporal information and not, yields the highest performance.
@@
Static target videos 25% 54.77 79.85 87.75 97.50
Dynamic target videos 75% 58.80 83.10 90.61 98.24
Nouns change 65% 58.06 82.43 89.98 97.97
Verbs change 09% 53.44 77.82 85.64 96.99
All 100% 59.82 83.84 91.28 98.24
D.5 Effect of modification text length
@@
(ours) 23.36 59.82 83.30 41.42 73.22 36.81 56.70
(longest) 33.35 58.84 82.72 35.64 67.60 33.20 52.13
This shows that while the generated modification texts are not as long on average as the ones from CIRR, our generated texts still provide useful training signal, as evidenced by the state-of-the-art performance of models trained on our dataset when transferred to standard CoIR tasks (CIRR and FashionIQ in Table IV and CIRCO in Table V). Moreover, the average number of characters of the modification text in WebVid-CoVR and the widely used FashionIQ CoIR benchmark are comparable (see Table I of the main paper: 23.36 vs 27.13, respectively).
D.6 Optimal number of frames
In the main paper, we sample 15 frames from WebVid-CoVR videos during training. Here, we experiment with incrementally increasing the number of frames for training and testing. We report the average recall on WebVid-CoVR-Test and observe a steady increase in performance with more frames (see Figure A.10). However, it’s worth noting that the performance is already quite high with just one frame, likely due to the static image bias inherent in the WebVid dataset, where videos are typically short in duration.
@@ D.6 Optimal number of frames
E Qualitative analysis